A 5.0 kg ball projected along the plane that is inclined 20° with the horizontal. What is its potential energy when it move 2.3 m up along that plane. Hint: use h = (length) (sinθ)
1
Expert's answer
2021-10-28T15:21:43-0400

"E_p=mgh=mgl\sin\theta=5\cdot 9.8\cdot 2.3\cdot \sin 20°=38.5~J."